# Single-step high-fidelity three-qubit gates by anisotropic chiral interactions

### v1

@@ I Introduction
Semiconductor spin qubits are a promising platform for quantum computing due to their long coherence times [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] and are rapidly emerging as a leading platform for scalable quantum computing due to their small footprint and compatibility with industrial semiconductor fabrication techniques [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Experimental advances have demonstrated single-qubit gate fidelities exceeding $99.9\%$ [26, 27] and two-qubit gate infidelities above $99\%$ in mid-scale devices [28, 29, 30, 31, 12]. Additionally, progress in coherent links enabling spin coupling over long distances holds promise for large-scale quantum processors with dense connectivity [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. With high-yield fabrication now achieved over large wafers [50], spin qubits are reaching the state where performing quantum computations is becoming increasingly viable [23, 51].
@@ II.1 Target gate
In this work, we focus on implementing in a single-step a doubly-controlled $\pi$-rotation by the $Y$-axis, i.e. a $\rm C^{2}Ry(\pi)$ gate, and we consider resonant gate protocols [58, 59, 12]. This gate is a non-Clifford operation, similar to the widely-used Toffoli gate; a fast and high-fidelity implementation of $\rm C^{2}Ry(\pi)$ enables a straightforward realization of the Toffoli gate without additional technical constraints, see e.g. the protocol proposed in Ref. [58]
@@
We illustrate the device in Fig. 1. This architecture is compatible with current silicon germanium based quantum processors [5, 62, 70, 64]. For concreteness, we now restrict ourselves to the $\rm C^{2}Ry(\pi)$ gate, where qubits $Q_{1}$ and $Q_{3}$ serve as control, and $Q_{2}$ is the target qubit. This means that $Q_{2}$ undergoes a $\rm Ry(\pi)$ rotation when $Q_{1}$ and $Q_{3}$ are in the state |$\ket{1}$. In matrix form, this gate reads
@@ II.2.2 Synchronization
Even in this ideal case, however, this gate implementation does not enable perfect fidelity. There are systematic errors originating from a close-in-frequency transitions. While driving the resonant transition |$\ket{101}\leftrightarrow\ket{111}$, we also inadvertently drive the other off-resonant transitions |$\ket{001}\leftrightarrow\ket{011}$, |$\ket{100}\leftrightarrow\ket{110}$, and |$\ket{000}\leftrightarrow\ket{010}$. These off-resonant transitions are the key limiting factors for achieving fast and high-fidelity three-qubit gates, as they require a low Rabi frequency $\Omega_{2}\ll J^{\parallel}_{12},J^{\parallel}_{23}$ or complex pulse shaping techniques to suppress them.
@@ II.4 Four-step echo-protocol
II.5 Systematic errors from flip-flop terms
@@
We classify the single-step gate protocol as follows. We refer to the gate as Iso and Ani gate if it is implemented using two-qubit exchange interactions, where the flip-flop terms take the form ${J_{ij}}(X_{i}X_{j}+Y_{i}Y_{j})/4$ (Iso gate) and ${J_{ij}}(X_{i}X_{j}-Y_{i}Y_{j})/4$ (Ani gate). The ChiralAni gate is an extension of the Ani gate that additionally incorporates the full three-qubit chiral anisotropic interaction discussed in the next section, including $ZZZ$ components. For gates implemented using the four-step protocol while utilizing the same interactions, we append (Echo) to the protocol name. We defer the discussion of the physical implementation of the interaction required for these gates to the next section, where we derive these interactions for spin qubit platforms.
@@
In this subsection, we have numerically simulated the gate protocols and shown that the inclusion of a $ZZZ$ interaction reduces the gate time needed to achieve a target fidelity and improve the gate fidelity for a given gate time. However, our analysis has thus far only considered the systematic errors arising from flip-flop interactions present in the Hamiltonian. In Section IV, we extend this analysis by performing numerical benchmarking of the various gate protocols under more realistic physical noise models, namely $1/f$ noise and calibration errors.
@@ III.1 Physical model
The flip-flop terms $\propto X_{i}X_{j}-Y_{i}Y_{j}$ (up to a single-qubit rotation around the $Z$-axis) couple the states {|$\{\ket{00}_{ij},\ket{11}_{ij}\}$ which have an energy gap $\omega_{i}+\omega_{j}$. As a result, the resonant gate condition requires the exchange interaction $J_{ij}$ to be small compared to $\omega_{i}+\omega_{j}$, which is usually an order of magnitude larger than $|\omega_{i}-\omega_{j}|$ in the isotropic exchange interaction, enabling faster controlled-rotation gates. The smallness of $J_{ij}$ compared to the energy gap $\omega_{i}+\omega_{j}$ also ensures that the flip-flop terms in Eq. (27) can be safely neglected when the gates are fast, justifying the ideal diagonal Hamiltonian introduced earlier in Eq. (2). We also numerically verify this assumption in Fig. 3, where the infidelity curve of the AniGate remains relatively constant as we vary the strength of the anisotropic flip-flop term by changing the gate time $T_{g}$.
@@ III.3.1 Isotropic chiral interaction
We now extend our effective spin model beyond the effective two-qubit interactions by considering higher-order terms in the expansion of Eq. (23). In a triangular dot arrangement, the third-order expansion introduces an effective three-qubit interaction, which we show to be non-negligible, with a magnitude on the order of $\sim\rm MHz$. In contrast, higher-order contributions, such as the fourth-order effective Hamiltonian $H_{\rm eff,4}$, primarily introduce minor corrections (on the order of $\sim 1~{}\rm kHz$) to the two-body interactions and are therefore neglected.
@@ III.3.2 Anisotropic chiral interaction with orbital magnetic field
For simplicity, we assume that all the SOI vectors $\vec{n}_{\rm so,ij}$ point in the $Z$-direction. In this case, as a spin tunnels from one dot to the next, it accumulates a certain spin-dependent phase because of SOI, in particular a spin-up (spin-down) particle accumulates a phase $e^{-i\gamma_{ij}}$ ($e^{i\gamma_{ij}}$) while tunneling from $Q_{i}$ to $Q_{j}$. This phase accumulation has an effect when the spin completes a closed loop around the triangular arrangement of QDs. An explicit example of two closed trajectories (neglecting the fermionic phase to simplify the notation) is
@@
The $ZZZ$ interaction arises from the destructive interference of these closed trajectories, where the spin-up and spin-down particles experience an accumulated phase-difference $\sum_{\langle ij\rangle}\gamma_{ij}$.
@@
We notice that the $ZZZ$ interaction is preceded by the prefactor $\sin(3\tilde{\gamma})$, which confirms that this interaction is indeed due to phase-interference effect between a spin-up particle and a spin-down particle after traversing the triangular loop. The effective Hamiltonian $\tilde{H}_{\rm eff,3}$ also contains single-qubit $Z$ operators and a modified scalar chirality term. The combination of the Pauli-$Z$ operators and the $ZZZ$ term ensures that there is no third-order virtual tunneling events if the spin-configuration is |$\ket{s,s,s}$. Physically, the Pauli-$Z$ operators produce small deviations to the qubit frequencies, which can be compensated for, and will be neglected in the following discussion. We also safely neglect the modified scalar chirality term (i.e in the upper-left block of Eq. (34b)) that is fully off-diagonal, and strongly suppressed by the large Zeeman energy difference, as we argued in Sec. II.
@@ III.4.2 Optimal setup
To suppress unwanted isotropic flip-flop terms between $Q_{1}$ and $Q_{3}$ caused by the exchange interaction $J_{13}$, we restrict $\tau_{13}/U$ to lie within the range $[0.02,0.05]$. Given our previous assumption that the distance between the QDs is $d_{QD}=100~{}\rm nm$, we estimate the required orbital magnetic field $B^{\perp}_{ext}$ to be in the range of $20~{}\rm mT$ to $60~{}\rm mT$. We also note that the required orbital magnetic field can be reduced by considering an alternative synchronization condition $(0,n,n,2n)$ with $n\geq 2$, see Eq. (53). The resulting gate will have lower gate speed but higher gate-fidelity due to the flip-flop terms scale inversely with $n$, see also Figs. 2 and 3. In the following, we show that the required orbital magnetic field is compatible with current silicon and germanium spin qubit devices.
@@ III.5 Measurement of the three-qubit interaction
We numerically verify that the 4-layer DD protocol can accurately detect the $ZZZ$ interaction with a high fidelity. The numerical performance of the more complex 32-layer DD protocol is provided in Appendix D. Here, we focus on the optimal SOI configuration for the $\rm C^{2}Ry(\pi)$ gate, presented in Fig. 4(c), with $\theta_{13}=0,\gamma=\pi/2$, considering an ideal system with all flip-flop interactions included.
@@ IV Gate performance with noise
Figure 6: Gate performance with noise Numerical simulation with qubits frequencies $\omega_{1}/2\pi=4.2~{}\rm GHz$, $\omega_{2}/2\pi=4.5~{}\rm GHz$, and $\omega_{3}/2\pi=4.8~{}\rm GHz$. Exchange interactions are $J_{12}/2\pi=J_{23}/2\pi=20~{}(90)~{}\rm MHz$ and $J_{13}/2\pi=1.6~{}\rm MHz$ for normal (fast) gates with gate time $T_{g}=100~{}(22.22)~{}\rm ns$. The Ani gate follows the $(0,1)$ synchronization, with Rabi frequencies $\Omega_{2}/2\pi\approx 5.16~{}(23.24)\rm MHz$, while the ChiralAni gate follows $(0,1,1,2)$, with $\Omega_{2}/2\pi\approx 5.04~{}(22.68)\rm MHz$. The three-qubit interaction amplitude is $J_{123}^{\parallel}/2\pi\approx 0.96~{}(4.34)~{}\rm MHz$. We plot the gate infidelity under different noise sources for the anisotropic interaction (red), chiral anisotropic interaction (blue), and their Echo variants (green and purple, respectively). (a-b) Gate infidelity against quasi-static error, where the noise is modeled as $\delta_{i}(t)=\delta_{i}$, drawn from a uniform distribution $[-\delta_{std},\delta_{std}]$. (c-d) Gate infidelity under high-frequency $1/f$ noise, modeled as a time-correlated process with power-density spectrum $PSD(f)=\delta_{std}/f$. Error bars represent $\pm$ one standard deviation from the mean.
@@
where we include two-qubit flip-flop terms and qubit-frequency shifts induced by the chiral anisotropic interaction, see Eq. (34).
@@
The noise terms $\delta_{i}(t)$ are treated as identically distributed random variables. For the quasistatic noise, we assume $\delta_{i}(t)$ is time-independent and drawn from a uniform distribution over the interval $[-\delta_{std},\delta_{std}]$. The choice of a uniform distribution is based on the assumption that noise is bounded by physical constraints, preventing rare but extreme fluctuations that can arise in a Gaussian model. This ensures that there is no unphysical outliers that could disproportionately affect the numerical average infidelity with small sample size. For the time-dependent noise model, $\delta(t)$ follows a time-correlated process with a power spectral density (PSD) given by $PSD(f)=\delta_{std}/f$. For both noise models, we numerically investigate up to $\delta_{std}=0.15$. For each value of $\delta_{std}$, we sample $200$ realizations. This noise model is similar to the model considered in Refs. [59, 58, 12, 114].
@@
For the numerical simulations presented in Fig. 6, we consider qubits with frequencies $\omega_{1}/2\pi=4.2~{}\rm GHz,$ $\omega_{2}/2\pi=4.5~{}\rm GHz,$ and $\omega_{3}/2\pi=4.8~{}\rm GHz$. The exchange interactions for the normal (fast) gate are $J_{12}/2\pi=J_{23}/2\pi=20~{}\rm MHz~{}(90~{}\rm MHz)$ and $J_{13}/2\pi=1.6~{}\rm MHz$. For the Ani gate, we adopt the synchronization condition $(0,1)$, resulting in a Rabi frequency of $\Omega_{2}/2\pi\approx 5.16~{}\rm MHz$ for the normal gate and $\Omega_{2}/2\pi\approx 23.24~{}\rm MHz$ for the fast gate. Meanwhile, for the ChiralAni gate, we use the synchronization condition $(0,1,1,2)$, yielding $\Omega_{2}/2\pi\approx 5.04~{}\rm MHz$ for the normal gate and $\Omega_{2}/2\pi\approx 22.68~{}\rm MHz$ for the fast gate. Correspondingly, the amplitude of the three-qubit interaction is $J_{123}^{\parallel}/2\pi\approx 0.96~{}\rm MHz$ for the normal gate and $J_{123}^{\parallel}/2\pi\approx 4.34~{}\rm MHz$ for the fast gate.
In Fig. 6, we observe that the fidelity of the ChiralAni gate decreases rapidly as we increase $\delta_{std}$ in both noise models and across both operating regimes. The only exception to this trend is found for the fast gate with $1/f$ noise in Fig. 6(c); this behavior is primarily due to the short gate time. Because of the short gate time, our numerical method only includes higher frequency components (from $\sim 10~{}\rm MHz$ up to $80~{}\rm GHz$) of the $1/f$ noise and does not account for the low-frequency components, which is partly captured by the quasi-static noise model. Nonetheless, the decrease in fidelity of the ChiralAni gate can be attributed entirely to the stochastic fluctuations of the two-qubit interactions. Given that the fidelity of the ChiralAni gate is already high ($1-\bar{F}\leq 10^{-4}$), even small amounts of noise will become dominant and significantly impact the fidelity.
We observe that the fidelity of the ChiralAni gate converges rapidly to the same as the one of the Ani gate, see Figs. 6(a) and 6(b). To investigate this further, we performed an identical numerical simulation with fluctuations only in the three-qubit interaction. In this case, the fidelity of the ChiralAni gate remains relatively stable up to $\delta_{std}=0.15$. Nevertheless, even with a moderate amount of noise $\delta=0.05$ (corresponding to fluctuations of up to $5\%$), the ChiralAni gate maintains an average fidelity of $\geq 99\%$.
@@
On the other hand, we observe in Fig. 6 that the four-step echo protocol (for both Ani and ChiralAni gates) is significantly more resilient to noise due to its echo structure. It typically outperforms the single-step protocol when moderate noise is present. A similar trend is seen in the ChiralAni (Echo) gate, where, due to its already high ideal fidelity, small two-qubit fluctuations become the dominant source of error. Consequently, the fidelity of the ChiralAni (Echo) gate converges to that of the Ani (Echo) gate, which is most clearly visible in Fig. 6(a). These results suggest that the four-step protocol could be well-suited for a wide range of near-term devices, where fluctuations of two-qubit interactions remain moderate. The single-step ChiralAni protocol remains the optimal choice if the noise level is small.

### v2

@@ I Introduction
Semiconductor spin qubits are a promising platform for quantum computing due to their long coherence times [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] and are rapidly emerging as a leading platform for scalable quantum computing due to their small footprint and compatibility with industrial semiconductor fabrication techniques [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Experimental advances have demonstrated single-qubit gate fidelities exceeding $99.9\%$ [26, 27] and two-qubit gate infidelities above $99\%$ in mid-scale devices [28, 29, 30, 31, 12, 32]. Additionally, progress in coherent links enabling spin coupling over long distances holds promise for large-scale quantum processors with dense connectivity [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. With high-yield fabrication now achieved over large wafers [51], spin qubits are reaching the state where performing quantum computations is becoming increasingly viable [23, 52].
@@ II.1 Target gate
In this work, we focus on implementing in a single-step a doubly-controlled $\pi$-rotation by the $Y$-axis, i.e. a $\rm C^{2}Ry(\pi)$ gate, and we consider resonant gate protocols [59, 60, 12]. This gate is a non-Clifford operation, similar to the widely-used Toffoli gate; a fast and high-fidelity implementation of $\rm C^{2}Ry(\pi)$ enables a straightforward realization of the Toffoli gate without additional technical constraints, see e.g. the protocol proposed in Ref. [59].
@@
We illustrate the device in Fig. 1. This architecture is compatible with current silicon germanium based quantum processors [5, 63, 71, 65]. The triple quantum dot system we study is a simple geometry that can exhibit ferromagnetism [72, 73, 74]. Its quantum transport properties have been extensively analyzed: the inherent $C_{3}$ symmetry gives rise to dark states that suppress particle current through the device [75, 76]. In addition, the $S=1/2$ sector of the triple quantum dot has been proposed as a decoherence-free subspace for encoding a logical qubit, offering built-in protection against global magnetic field fluctuations [77, 78, 79]. For a thorough review of the triple quantum dot physics, we refer to Ref. [80].
For concreteness, we now restrict ourselves to the $\rm C^{2}Ry(\pi)$ gate, where qubits $Q_{1}$ and $Q_{3}$ serve as control, and $Q_{2}$ is the target qubit. This means that $Q_{2}$ undergoes a $\rm Ry(\pi)$ rotation when $Q_{1}$ and $Q_{3}$ are in the state |$\ket{1}$. In matrix form, this gate reads
@@ II.2.2 Synchronization
Even in this ideal case, however, this gate implementation does not enable perfect fidelity. There are systematic errors originating from a close-in-frequency transitions. While driving the resonant transition |$\ket{101}\leftrightarrow\ket{111}$, we also inadvertently drive the other off-resonant transitions |$\ket{001}\leftrightarrow\ket{011}$, |$\ket{100}\leftrightarrow\ket{110}$, and |$\ket{000}\leftrightarrow\ket{010}$. These off-resonant transitions are the key limiting factors for achieving fast and high-fidelity three-qubit gates, as they require a low Rabi frequency $\Omega_{2}\ll J^{\parallel}_{12},J^{\parallel}_{23}$ or complex pulse shaping techniques [81, 82, 83, 84] to suppress them.
@@ II.3 Ideal gate with three-qubit interaction
In this section, we demonstrated that introducing a small three-qubit $ZZZ$ interaction can resolve the synchronization issue in the implementation of the $\rm C^{2}Ry(\pi)$ gate. However, we emphasize that this approach is not limited to this specific gate. As shown in Ref. [87], an ideal $\rm C^{2}Z$ gate can also be realized using an additional $ZZZ$ interactions. More broadly, such three-body interactions enable the implementation of a wide class of three-qubit gates that act non-trivially on a designated two-dimensional subspace of the full eight-dimensional Hilbert space, while leaving the orthogonal subspace invariant.
@@ II.4 Four-step echo-protocol
II.5 Systematic errors from flip-flop and flop-flop terms
@@
We classify the single-step gate protocol as follows. We refer to the gate as Iso and Ani gate if it is implemented using two-qubit exchange interactions, where we have flip-flop terms ${J_{ij}}(X_{i}X_{j}+Y_{i}Y_{j})/4$ for the Iso gate and flop-flop terms ${J_{ij}}(X_{i}X_{j}-Y_{i}Y_{j})/4$ for the Ani gate. The ChiralAni gate is an extension of the Ani gate that additionally incorporates the full three-qubit chiral anisotropic interaction discussed in the next section, including $ZZZ$ components. For gates implemented using the four-step protocol while utilizing the same interactions, we append (Echo) to the protocol name. We defer the discussion of the physical implementation of the interaction required for these gates to the next section, where we derive these interactions for spin qubit platforms.
@@
$J_{12}$ $J_{23}$ $J_{13}$ $J_{123}$ Echo $1-\bar{F}$ $30~{}\mathrm{ns}$ $1-\bar{F}$ $100~{}\mathrm{ns}$
Fi-Fo Fi-Fo N/A N/A $\times$ $\approx 0.24$ $\approx 0.04$
Fo-Fo Fo-Fo N/A N/A $\times$ $\approx 0.023$ $\approx 0.02$
Fo-Fo Fo-Fo N/A N/A $\checkmark$ $\approx 7\times 10^{-4}$ $\approx 4\times 10^{-4}$
Fo-Fo Fo-Fo Fi-Fo ChiralAni $\times$ $\approx 7\times 10^{-4}$ $\approx 6\times 10^{-5}$
Fo-Fo Fo-Fo Fi-Fo ChiralAni $\checkmark$ $\approx 3\times 10^{-4}$ $\approx 4\times 10^{-5}$
Table 1: Summary of gate protocol performance under flip-flop and flop-flop interactions. Each protocol is defined by the off-diagonal elements of the pairwise exchange terms $J_{ij}$, either flip-flop (Fi-Fo: $X_{i}X_{j}+Y_{i}Y_{j}$) or flop-flop (Fo-Fo: $X_{i}X_{j}-Y_{i}Y_{j}$), as well as the presence of a chiral three-qubit interaction $J_{123}$ and the use of the proposed echo sequence. The final two columns report simulated average gate infidelity $1-\bar{F}$ at gate durations $T_{g}=30~{}\mathrm{ns}$ and $T_{g}=100~{}\mathrm{ns}$.
@@
In this subsection, we have numerically simulated the gate protocols and shown that the inclusion of a $ZZZ$ interaction reduces the gate time needed to achieve a target fidelity and improve the gate fidelity for a given gate time. For clarity, Table 1 summarizes the five gate protocols shown in Fig. 3, including the required two- and three-qubit interactions and the average gate infidelities at $T_{g}=30~{}\rm ns$ and $T_{g}=\rm 100~{}\rm ns$. However, our analysis has thus far only considered the systematic errors arising from flip-flop/flop-flop interactions present in the Hamiltonian. In Section IV, we extend this analysis by performing numerical benchmarking of the various gate protocols under more realistic physical noise models, namely $1/f$ noise and calibration errors.
@@ III.1 Physical model
The flop-flop terms $\propto X_{i}X_{j}-Y_{i}Y_{j}$ (up to a single-qubit rotation around the $Z$-axis) couple the states {|$\{\ket{00}_{ij},\ket{11}_{ij}\}$ which have an energy gap $\omega_{i}+\omega_{j}$. As a result, the resonant gate condition requires the exchange interaction $J_{ij}$ to be small compared to $\omega_{i}+\omega_{j}$, which is usually an order of magnitude larger than $|\omega_{i}-\omega_{j}|$ in the isotropic exchange interaction, enabling faster controlled-rotation gates. The smallness of $J_{ij}$ compared to the energy gap $\omega_{i}+\omega_{j}$ also ensures that the flop-flop terms in Eq. (27) can be safely neglected when the gates are fast, justifying the ideal diagonal Hamiltonian introduced earlier in Eq. (2). We also numerically verify this assumption in Fig. 3, where the infidelity curve of the Ani gate remains relatively constant as we vary the strength of the anisotropic flop-flop term by changing the gate time $T_{g}$.
@@ III.3.1 Isotropic chiral interaction
We now extend our effective spin model beyond the effective two-qubit interactions by considering higher-order terms in the expansion of Eq. (23). In a triangular dot arrangement, the third-order expansion introduces an effective three-qubit interaction, which we show to be non-negligible, with a magnitude on the order of $\sim\rm MHz$. In contrast, higher-order contributions, such as the fourth-order effective Hamiltonian $H_{\rm eff,4}$, primarily introduce minor corrections (on the order of $\sim 1~{}\rm kHz$) to the two-body interactions [95] and are therefore neglected.
@@ III.3.2 Anisotropic chiral interaction with orbital magnetic field
Because of the fermionic anticommutation relations, our system presents two distinct destructive interference effects. The $ZZZ$ interaction emerges from the intricate interplay of both destructive interference effects.
The first comes from interference between a trajectory and its time-reversed version. Without an orbital magnetic field, chiral symmetry ensures $[\tau_{ij}]_{\sigma\sigma}=[\tau_{ji}]_{\bar{\sigma}\bar{\sigma}}$. With an orbital magnetic field, each trajectory accumulates a total magnetic phase $\Phi_{B}$ while traveling clockwise and $-\Phi_{B}$ counterclockwise. For example, the accumulated phases along a trajectory and its time-reversed counterpart are
@@
As a result, the longitudinal interaction $J_{123}^{\parallel}$ acquires the factor $\sin(\Phi_{B})$ because of the Aharonov-Bohm effect.
The second interference effect arises from the spin-dependent phase difference induced by SOI. For simplicity, we assume that all SOI vectors $\vec{n}_{\rm so,ij}$ point in the $Z$-direction. In this case, as a spin tunnels from one dot to the next, it accumulates a certain spin-dependent phase because of SOI, in particular a spin-up (spin-down) particle accumulates a phase $e^{-i\gamma_{ij}}$ ($e^{i\gamma_{ij}}$) while tunneling from $Q_{i}$ to $Q_{j}$. This phase accumulation has an effect when the spin completes a closed loop around the triangular arrangement of QDs. An example of two closed trajectories with the same chirality (i.e., the same Aharonov-Bohm phase) is
[…]
Consequently, the longitudinal interaction $J_{123}^{\parallel}$ acquires a factor of $\sin(\sum_{\langle ij\rangle}\gamma_{ij})$ due to the SOI-induced phase difference between spin-up and spin-down particles. In Appendix A, we explicitly derive the dependence of the longitudinal interaction $J_{123}^{\parallel}$ on third-order virtual tunneling amplitudes, see e.g. Eq. (63).
@@
We notice that the $ZZZ$ interaction is preceded by the prefactor $\sin(3\tilde{\gamma})$, which confirms that this interaction is indeed due to the phase-interference effect between a spin-up particle and a spin-down particle after traversing the triangular loop. We also find that the effective Hamiltonian $\tilde{H}_{\rm eff,3}$ contains the factor $\sin(\Phi_{B})$ due to chirality symmetry breaking. The effective Hamiltonian $\tilde{H}_{\rm eff,3}$ also contains single-qubit $Z$ operators and a modified scalar chirality term. The combination of the Pauli-$Z$ operators and the $ZZZ$ term ensures that there is no third-order virtual tunneling events if the spin-configuration is |$\ket{s,s,s}$. Physically, the Pauli-$Z$ operators produce small deviations to the qubit frequencies, which can be compensated for, and will be neglected in the following discussion. We also safely neglect the modified scalar chirality term (i.e in the upper-left block of Eq. (35b)) that is fully off-diagonal, and strongly suppressed by the large Zeeman energy difference, as we argued in Sec. II.
@@ III.4.2 Optimal setup
To suppress unwanted flip-flop terms between $Q_{1}$ and $Q_{3}$ caused by the exchange interaction $J_{13}$, we restrict $\tau_{13}/U$ to lie within the range $[0.02,0.05]$. Given our previous assumption that the distance between the QDs is $d_{QD}=100~{}\rm nm$, we estimate the required orbital magnetic field $B^{\perp}_{ext}$ to be in the range of $20~{}\rm mT$ to $60~{}\rm mT$. We also note that the required orbital magnetic field can be reduced by considering an alternative synchronization condition $(0,n,n,2n)$ with $n\geq 2$, see Eq. (55). The resulting gate will have lower gate speed but higher gate-fidelity due to the flip-flop terms scale inversely with $n$, see also Figs. 2 and 3. In the following, we show that the required orbital magnetic field is compatible with current silicon and germanium spin qubit devices.
@@ III.5 Measurement of the three-qubit interaction
We numerically verify that the 4-layer DD protocol can accurately detect the $ZZZ$ interaction with a high fidelity. The numerical performance of the more complex 32-layer DD protocol is provided in Appendix D. Here, we focus on the optimal SOI configuration for the $\rm C^{2}Ry(\pi)$ gate, presented in Fig. 4(c), with $\theta_{13}=0,\gamma=\pi/2$, considering an ideal system with all flip-flop/flop-flop interactions included.
@@ IV Gate performance with noise
Figure 6: Gate performance with noise Numerical simulation with qubits frequencies $\omega_{1}/2\pi=4.2~{}\rm GHz$, $\omega_{2}/2\pi=4.5~{}\rm GHz$, and $\omega_{3}/2\pi=4.8~{}\rm GHz$. Exchange interactions are $J_{12}/2\pi=J_{23}/2\pi=20~{}(90)~{}\rm MHz$ and $J_{13}/2\pi=1.6~{}\rm MHz$ for normal (fast) gates with gate time $T_{g}=100~{}(22.22)~{}\rm ns$. The Ani gate follows the $(0,1)$ synchronization, with Rabi frequencies $\Omega_{2}/2\pi\approx 5.16~{}(23.24)\rm MHz$, while the ChiralAni gate follows $(0,1,1,2)$, with $\Omega_{2}/2\pi\approx 5.04~{}(22.68)\rm MHz$. The three-qubit interaction amplitude is $J_{123}^{\parallel}/2\pi\approx 0.96~{}(4.34)~{}\rm MHz$. We plot the gate infidelity under different noise sources for the anisotropic interaction (red), chiral anisotropic interaction (blue), and their Echo variants (green and purple, respectively). (a-b) Gate infidelity against quasi-static error, where the noise is modeled as $\delta_{i}(t)=\delta_{i}$, drawn from a uniform distribution $[-\delta_{std},\delta_{std}]$. (c-d) Gate infidelity under high-frequency $1/f$ noise, modeled as a white noise with fluctuation $\delta_{std}$ under infinite impulse response filter. Error bars represent $\pm$ one standard deviation from the mean.
@@
where we include two-qubit flip-flop/flop-flop terms and qubit-frequency shifts induced by the chiral anisotropic interaction, see Eq. (35).
Physically, our noise model accounts for low-frequency fluctuations in tunneling amplitudes induced by variations in the barrier gate voltages $V_{ij}$. These fluctuations cause variations in both two-body exchange and three-body interaction terms, which can be comparable in magnitude relative to their respective strengths. Physically, these fluctuations can be attributed to charge noise, a critical noise source in semiconductor spin qubits [8]. We adopt this model because tunnel coupling fluctuations are particularly detrimental: they shift the resonance frequencies Eq. (4) and affect the synchronization ratios Eqs. (17). Although more realistic noise models could be considered, we believe that the one employed here effectively captures the essential features that influence gate performance under realistic conditions.
@@
The noise terms $\delta_{i}(t)$ are treated as identically distributed random variables. For the quasistatic error, we assume $\delta_{i}(t)$ is time-independent, and drawn from a uniform distribution over the interval $[-\delta_{std},\delta_{std}]$. The choice of a uniform distribution is based on the assumption that noise is bounded by physical constraints, preventing rare but extreme fluctuations that can arise in a Gaussian model. This ensures that there is no unphysical outliers that could disproportionately affect the numerical average infidelity with small sample size.
For the time-dependent noise model, $\delta(t)$ follows a time-correlated process with a power spectral density (PSD) given by $PSD(f)\propto 1/f$. We generate the $1/f$ noise with the same order of fluctuation $\delta_{std}$ as a white noise using the infinite impulse response filter (with 300 poles) described in Ref. [128]. We normalize the discrete $1/f$ noise such that on average it has the same power as white noises with fluctuation $\delta_{std}$
$\frac{\sum_{n=0}^{N_{\rm samples}}PSD(nf_{s})}{N_{\rm samples}}=2\delta_{std}^{2}\Delta t.$ (51)
Because of the short gate time, our numerical method only includes high-frequency components (from $\sim 2~{}\rm MHz$ up to $80~{}\rm GHz$) of the $1/f$ noise. The low-frequency components is partly captured by the quasi-static noise model.
For both noise models, we investigate numerically up to $\delta_{std}=0.15$ with discrete step $\Delta\delta_{std}=0.03$. For each value of $\delta_{std}$, we sample $200$ realizations. This noise model is similar to the model considered in Refs. [60, 59, 12, 129].
@@
For the numerical simulations presented in Fig. 6, we consider qubits with frequencies $\omega_{1}/2\pi=4.2~{}\rm GHz,$ $\omega_{2}/2\pi=4.5~{}\rm GHz,$ and $\omega_{3}/2\pi=4.8~{}\rm GHz$. The exchange interactions for the normal (fast) gate are $J_{12}/2\pi=J_{23}/2\pi=20~{}\rm MHz~{}(90~{}\rm MHz)$ and $J_{13}/2\pi=1.6~{}\rm MHz$. For the Ani gate, we adopt the synchronization condition $(0,1)$, resulting in a Rabi frequency of $\Omega_{2}/2\pi\approx 5.16~{}\rm MHz$ for the normal gate and $\Omega_{2}/2\pi\approx 23.24~{}\rm MHz$ for the fast gate. Meanwhile, for the ChiralAni gate, we use the synchronization condition $(0,1,1,2)$, yielding $\Omega_{2}/2\pi\approx 5.04~{}\rm MHz$ for the normal gate and $\Omega_{2}/2\pi\approx 22.68~{}\rm MHz$ for the fast gate. Correspondingly, the amplitude of the three-qubit interaction is $J_{123}^{\parallel}/2\pi\approx 0.96~{}\rm MHz$ for the normal gate and $J_{123}^{\parallel}/2\pi\approx 4.34~{}\rm MHz$ for the fast gate. In Figs. 6(a) and 6(b), we observe that the ChiralAni gate fidelity decreases rapidly with increasing $\delta_{std}$ in both operating regimes but has similar performance as the other gates when considering $1/f$ noise as shown in Figs. 6(c) and 6(d). However, we note that due to the discrete nature of our simulation, low-frequency components of $1/f$ noise manifest as quasi-static errors. Consequently, we find the ChiralAni gate fidelity rapidly converges to that of the Ani gate (Figs. 6(a) and 6(b)).
To investigate further, we conducted simulations where only the three-qubit interaction was subject to noise, revealing that the ChiralAni gate remains stable up to $\delta_{std}=0.15$. Therefore, we conclude that the fidelity drop is entirely due to stochastic fluctuations in two-qubit interactions. Since the ChiralAni gate already achieves high fidelity ($1-\bar{F}\leq 10^{-4}$) without noise, even small noise contributions become dominant and significantly impact performance. Nevertheless, even with a moderate amount of noise $\delta=0.05$ (corresponding to $5\%$ fluctuations), the ChiralAni gate maintains an average fidelity of $\geq 99\%$.
@@
On the other hand, we observe in Fig. 6 that the four-step echo protocol (for both Ani and ChiralAni gates) is significantly more resilient to noise due to its echo structure. It typically outperforms the single-step protocol when moderate noise is present. A similar trend is seen in the ChiralAni (Echo) gate, where, due to its already high ideal fidelity, small two-qubit fluctuations become the dominant source of error. Consequently, the fidelity of the ChiralAni (Echo) gate converges to that of the Ani (Echo) gate. This trend is clearly visible for example in Fig. 6(a). These results suggest that the four-step protocol could be well-suited for a wide range of near-term devices, where fluctuations of two-qubit interactions remain moderate. The single-step ChiralAni protocol remains the optimal choice if the noise level is small.